OpenStudy (anonymous):

If F(x) = x² + 5x and G(x) = 2x + 1, find F(5) + G(6)

OpenStudy (anonymous):

for F(5), replace x in the F(x) = x² + 5x with 5... and for G(6) replace x in the G(x) = 2x + 1 with 6... and then add them you'll get the answer.. Got it..??

OpenStudy (anonymous):

f(5)=5^5+5^5=50 g(6)=2^6+1=13 f(5)+g(6)=50+13=63

OpenStudy (anonymous):

the steps of upper solution are wrong.. but answer is right... lol

OpenStudy (anonymous):

@ryantorres got it??

OpenStudy (anonymous):

yeah , i got it . really helpful

OpenStudy (anonymous):

so you should write it... (2)(6) or 2x6 ....